OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 4x – 3y = –14 4x + y = –14 6x + 5y = 13 6x – 5y = –13

OpenStudy (anonymous):

simplify the last equation then add them... what do you get? 2x – y + 2z = –15 2*(2x – 2y – z = 1)

OpenStudy (anonymous):

2x – y + 2z = –15 4x –4y – 2z = 2 add these two now...